# Non-Convex Stochastic Composite Optimization with Polyak Momentum

### v1

@@ Appendix A Missing proofs in Section 5
We put together Lemma 5.2 and Lemma 5.3 and get:
$\Phi_{k+1}$ $=F_{k+1}+a\Delta_{k+1}$
$\leq F_{k}-\frac{M-L}{4}R_{k}+\frac{\Delta_{k}}{M-L}$
$\quad+a\left[(1-\gamma)\Delta_{k}+\frac{L^{2}}{\gamma}R_{k}+\gamma^{2}\sigma^{2}\right]$
$=F_{k}-\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)R_{k}+a\gamma^{2}\sigma^{2}$
$\quad+\left(1-\gamma+\frac{1}{a(M-L)}\right)a\Delta_{k}$
If $\frac{M-L}{4}-\frac{aL^{2}}{\gamma}>0$, then we can plug Lemma 5.5 in and get:
$\Phi_{k+1}$ $\leq F_{k}-\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)(\frac{{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{3(M^{2}+L^{2})}-\frac{\Delta_{k}}{M^{2}+L^{2}})$
$\quad+\left(1-\gamma+\frac{1}{a(M-L)}\right)a\Delta_{k}+a\gamma^{2}\sigma^{2}$
$=F_{k}-\left(\frac{M-L}{12(M^{2}+L^{2})}-\frac{aL^{2}}{3\gamma(M^{2}+L^{2})}\right){\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]$
$\quad+\left(1-\gamma+\frac{1}{a(M-L)}+\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)\frac{1}{a(M^{2}+L^{2})}\right)a\Delta_{k}+a\gamma^{2}\sigma^{2}$
Now let
$\frac{M-L}{24(M^{2}+L^{2})}=\frac{aL^{2}}{3\gamma(M^{2}+L^{2})}$
then $a=\frac{\gamma(M-L)}{8L^{2}}$. Note that with this $a$, we have $\frac{M-L}{4}-\frac{aL^{2}}{\gamma}>0$. Now
$\frac{1}{a(M-L)}+\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)\frac{1}{a(M^{2}+L^{2})}$ $=\frac{8L^{2}}{\gamma(M-L)^{2}}+\frac{M-L}{8}\frac{1}{a(M^{2}+L^{2})}$
$=\frac{8L^{2}}{\gamma(M-L)^{2}}+\frac{L^{2}}{\gamma(M^{2}+L^{2})}$
$=\frac{L^{2}}{\gamma}\left(\frac{8}{(M-L)^{2}}+\frac{1}{M^{2}+L^{2}}\right)$
Therefore we need $\frac{L^{2}}{\gamma}\left(\frac{8}{(M-L)^{2}}+\frac{1}{M^{2}+L^{2}}\right)\leq\gamma$, rearranging, it suffices to have:
$\gamma^{2}\geq L^{2}\left(\frac{8}{(M-L)^{2}}+\frac{1}{M^{2}+L^{2}}\right)$
Note that $(M-L)^{2}\leq M^{2}+L^{2}$, it suffices to set $\gamma^{2}=\frac{9L^{2}}{(M-L)^{2}}$, or equivalently:
$\gamma=\frac{3L}{(M-L)}$
Note that this requires $M>4L$. Now
$a=\frac{M-L}{8L^{2}}\frac{3L}{M-L}=\frac{3}{8L}$
Putting them all together,
$\Phi_{k+1}\leq\Phi_{k}-\frac{M-L}{24(M^{2}+L^{2})}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{27L}{8(M-L)^{2}}\sigma^{2}$
Let $M=\tau L$ for some $\tau>4$, then
$\Phi_{k+1}\leq\Phi_{k}-\frac{\tau-1}{24(\tau^{2}+1)L}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{27}{8(\tau-1)^{2}L}\sigma^{2}$
Note that for $\tau>4,\frac{\tau-1}{\tau^{2}+1}\geq\frac{1}{2\tau}$ and $\frac{1}{(\tau-1)^{2}}\leq\frac{2}{\tau^{2}}$. Therefore, for $M>(1+3\sqrt{3})L$:
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{48M}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{27L}{4M^{2}}\sigma^{2}$
Rearranging and summing Equation 7 from 00 to $K-1$, we get
@@
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\mathcal{O}(\sqrt{\frac{L\Phi_{0}\sigma^{2}}{K}}+\frac{L\Phi_{0}}{K})$
@@
Appendix B Missing proofs in Section 6
See 6.1
@@ Appendix B Missing proofs in Section 6
In the proof of Lemma 5.6, if we set
$\frac{L^{2}}{\gamma}\left(\frac{8}{(M-L)^{2}}+\frac{1}{M^{2}+L^{2}}\right)\leq\frac{\gamma}{2},$
or equivalently,
$\gamma=\frac{3\sqrt{2}L}{M-L},$
Note that this requires $M>(1+3\sqrt{2})L$. Now we have $a=\frac{\sqrt{2}}{8L}$ and:
$\Phi_{k+1}\leq\Phi_{k}-\frac{M-L}{24(M^{2}+L^{2})}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]-\frac{3\sqrt{2}L}{2(M-L)}\Delta_{k}+\frac{27\sqrt{2}L}{4(M-L)^{2}}\sigma^{2}$
Let $M=\tau L$ for some $\tau>(1+3\sqrt{2})$, then:
$\Phi_{k+1}\leq\Phi_{k}-\frac{\tau-1}{24(\tau^{2}+1)L}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]-\frac{3\sqrt{2}}{2(\tau-1)}\Delta_{k}+\frac{27\sqrt{2}}{4(\tau-1)^{2}L}\sigma^{2}$
Note that for $\tau>(1+3\sqrt{2}),\frac{\tau-1}{\tau^{2}+1}\geq\frac{1}{2\tau},\frac{1}{2(\tau-1)}\geq\frac{1}{2\tau}$ and $\frac{1}{(\tau-1)^{2}}\leq\frac{2}{\tau^{2}}$. Therefore, for $M>(1+3\sqrt{2})L$:
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{48M}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]-\frac{3\sqrt{2}}{2M}\Delta_{k}+\frac{27\sqrt{2}L}{4M^{2}}\sigma^{2}$
@@ Appendix C Missing proofs in Section 7
First, we notice that under 3.1 and 3.3, Lemma 5.3 still holds in the inexact case. Now we give an analogous result to Lemma 5.5 in the inexact case:
Lemma C.1.
Given 3.3, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, we have:
$R_{k}\geq\frac{{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{4(M^{2}+L^{2})}-\frac{\Delta_{k}}{M^{2}+L^{2}}-\frac{{\mathbb{E}}\left[\left\lVert\nabla\Omega(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{M^{2}+L^{2}}$
@@
$=\left\lVert M(\mathbf{x}_{k}-\mathbf{x}_{k+1})+\nabla\Omega(\mathbf{x}_{k+1})+(\nabla f(\mathbf{x}_{k})-\mathbf{m}_{k})+(\nabla f(\mathbf{x}_{k+1})-\nabla f(\mathbf{x}_{k}))\right\rVert^{2}$
$\leq 4M^{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+4\left\lVert\nabla\Omega(\mathbf{x}_{k+1})\right\rVert^{2}+4\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}+4\left\lVert\nabla f(\mathbf{x}_{k+1})-\nabla f(\mathbf{x}_{k})\right\rVert^{2}$
$\leq 4(M^{2}+L^{2})\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+4\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}+4\left\lVert\nabla\Omega(\mathbf{x}_{k+1})\right\rVert^{2}$
Rearranging, we have:
$-\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}\leq-\frac{\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}}{4(M^{2}+L^{2})}+\frac{\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}}{M^{2}+L^{2}}+\frac{\left\lVert\nabla\Omega(\mathbf{x}_{k+1})\right\rVert^{2}}{M^{2}+L^{2}}$
Notice that Lemma 5.2 still holds. We continue to use the same Lyapunov function $\Phi_{k}\coloneqq F_{k}+a\Delta_{k}$.
Lemma C.2.
Given 3.1, 3.3, and 3.4, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, if $\gamma\coloneqq\sqrt{\frac{456}{17}}\frac{L}{M-L},a\coloneqq\sqrt{\frac{19}{408}}\frac{1}{L}$, that in each iteration we have:
$\left\lVert\nabla\Omega(\mathbf{x}_{k+1})\right\rVert^{2}\leq\frac{M^{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}}{16}+S_{k},$
then if $M>7L$, we have:
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{68M}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{12L}{M^{2}}\sigma^{2}+\frac{2}{17M}S_{k}$ (10)
@@
First we plug Equation 9 into Lemma C.1 and get:
$-R_{k}$ $\leq-\frac{{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{4(M^{2}+L^{2})}+\frac{\Delta_{k}}{M^{2}+L^{2}}+\frac{{\mathbb{E}}\left[\left\lVert\nabla\Omega(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{M^{2}+L^{2}}$
$\leq-\frac{{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{4(M^{2}+L^{2})}+\frac{\Delta_{k}}{M^{2}+L^{2}}+\frac{R_{k}}{16}+\frac{S_{k}}{M^{2}+L^{2}}$

### v4

@@ Appendix B Missing Proofs in Section 5
We put together Lemmas 1 and 2 and get:
$\Phi_{k+1}$ $=F_{k+1}+a\Delta_{k+1}\leq F_{k}-\frac{M_{k}-L}{4}R_{k}+\frac{\Delta_{k}}{M_{k}-L}+a\Bigl{[}(1-\gamma_{k})\Delta_{k}+\frac{L^{2}}{\gamma_{k}}R_{k}+\gamma_{k}^{2}\sigma^{2}\Bigr{]}$
$=F_{k}-H_{k}R_{k}+a\gamma_{k}^{2}\sigma^{2}+\left(1-\gamma_{k}+\frac{1}{a(M_{k}-L)}\right)a\Delta_{k},$
where $H_{k}\coloneqq\frac{M_{k}-L}{4}-\frac{aL^{2}}{\gamma_{k}}$. If $H_{k}>0$, then we can plug Lemma 4 in and get:
$\Phi_{k+1}$ $\leq F_{k}-\frac{H_{k}}{M_{k}^{2}+L^{2}}\left(\frac{1}{3}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]-\Delta_{k}\right)+\left(1-\gamma_{k}+\frac{1}{a(M_{k}-L)}\right)a\Delta_{k}+a\gamma_{k}^{2}\sigma^{2}$
$=F_{k}-\frac{H_{k}}{3(M_{k}^{2}+L^{2})}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\left(1-\gamma_{k}+\frac{1}{a(M_{k}-L)}+\frac{H_{k}}{a(M_{k}^{2}+L^{2})}\right)a\Delta_{k}+a\gamma_{k}^{2}\sigma^{2}.$
Now let us choose $a=\frac{\gamma_{k}(M_{k}-L)}{8L^{2}}$, so that $H_{k}=\frac{M_{k}-L}{8}>0$. Then,
$\frac{1}{a(M_{k}-L)}+\frac{H_{k}}{a(M_{k}^{2}+L^{2})}=\frac{8L^{2}}{\gamma_{k}(M_{k}-L)^{2}}+\frac{L^{2}}{\gamma_{k}(M_{k}^{2}+L^{2})}=\frac{L^{2}}{\gamma_{k}}\left(\frac{8}{(M_{k}-L)^{2}}+\frac{1}{M_{k}^{2}+L^{2}}\right).$
Therefore we need $\frac{L^{2}}{\gamma_{k}}\bigl{(}\frac{8}{(M_{k}-L)^{2}}+\frac{1}{M_{k}^{2}+L^{2}}\bigr{)}\leq\gamma_{k}$ or, equivalently, $\gamma_{k}^{2}\geq L^{2}\bigl{(}\frac{8}{(M_{k}-L)^{2}}+\frac{1}{M_{k}^{2}+L^{2}}\bigr{)}$. Since $(M_{k}-L)^{2}\leq M_{k}^{2}+L^{2}$, it suffices to set $\gamma_{k}^{2}=\frac{9L^{2}}{(M_{k}-L)^{2}}$, i.e., $\gamma_{k}=\frac{3L}{M_{k}-L}$. Note that this requires $M_{k}>4L$ as we want to keep $\gamma_{k}<1$. For our value of $\gamma_{k}$, we get $a=\frac{3}{8L}$. Putting everything together, we obtain
$\Phi_{k+1}\leq\Phi_{k}-\frac{M_{k}-L}{24(M_{k}^{2}+L^{2})}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{27L\sigma^{2}}{8(M_{k}-L)^{2}}.$
Since $\frac{M_{k}-L}{M_{k}^{2}+L^{2}}\geq\frac{1}{2(M_{k}-L)}\geq\frac{1}{2M_{k}}$, we can further estimate
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{48M_{k}}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{27L\sigma^{2}}{4M_{k}^{2}}.\qed$
See 6
Proof.
Setting $M_{k}\coloneqq M$ some constant, rearranging and summing Equation 7 from 00 to $K-1$, we get
@@
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\frac{192L\Phi_{0}}{K}+48(3^{\nicefrac{{3}}{{2}}})\sqrt{\frac{L\Phi_{0}\sigma^{2}}{K}}$
@@
If $t$ is chosen uniformly at random from $\{1,\ldots,K\}$, then ${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{t})\right\rVert^{2}\right]=\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]$. ∎
We also discuss the case where $M_{k}$ is not a constant: See 7
@@
Denote $G_{k+1}^{2}\coloneqq{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]$. According to Lemma 5, we have
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{48M_{k}}G_{k+1}^{2}+\frac{27L\sigma^{2}}{4M_{k}^{2}}.$
Rearranging and summing the above from 00 to $K-1$, where $K\geq 1$ is arbitrary, we get:
$\frac{1}{48}\sum_{k=0}^{K-1}\frac{1}{M_{k}}G_{k+1}^{2}\leq\Phi_{0}+\frac{27L\sigma^{2}}{4}\sum_{k=0}^{K-1}\frac{1}{M_{k}^{2}}.$
Denoting $A_{i}=\sum_{k=0}^{i-1}\frac{1}{M_{k}}$, we obtain
$\frac{1}{A_{K}}\sum_{k=0}^{K-1}\frac{1}{M_{k}}G_{k+1}^{2}\leq\frac{48\Phi_{0}+324L\sigma^{2}\sum_{k=0}^{K-1}\frac{1}{M_{k}^{2}}}{A_{K}},$
Hence, for $M_{k}\coloneqq\max\bigl{\{}\sqrt{\frac{(k+1)L\sigma^{2}}{\phi_{0}}},4L\bigr{\}}$, we have
$A_{K}=\sum_{k=1}^{K}\min\biggl{\{}\sqrt{\frac{\Phi_{0}}{kL\sigma^{2}}},\frac{1}{4L}\biggr{\}}\geq K\min\biggl{\{}\sqrt{\frac{\Phi_{0}}{KL\sigma^{2}}},\frac{1}{4L}\biggr{\}}=\min\biggl{\{}\sqrt{\frac{K\Phi_{0}}{L\sigma^{2}}},\frac{K}{4L}\biggr{\}}$
and
$48\Phi_{0}+324L\sigma^{2}\sum_{k=0}^{K-1}\frac{1}{M_{k}^{2}}\leq 48\Phi_{0}+324L\sigma^{2}\sum_{k=1}^{K}\frac{\Phi_{0}}{kL\sigma^{2}}\leq 372\Phi_{0}\sum_{k=1}^{K}\frac{1}{k}.$
Putting everything together, we get:
$\frac{1}{A_{K}}\sum_{k=0}^{K-1}\frac{1}{M_{k}}G_{k+1}^{2}\leq\frac{372\Phi_{0}\sum_{k=1}^{K-1}\frac{1}{k}}{\min\{\sqrt{\frac{K\Phi_{0}}{L\sigma^{2}}},\frac{K}{4L}\}}\leq\frac{372\Phi_{0}(\ln K+1)}{\min\{\sqrt{\frac{K\Phi_{0}}{L\sigma^{2}}},\frac{K}{4L}\}}\leq 372(\ln K+1)\Bigl{(}\sqrt{\frac{L\Phi_{0}\sigma^{2}}{K}}+\frac{L\Phi_{0}}{K}\Bigr{)}.$
If $t(k)$ is chosen from $\{1,\ldots,k\}$ with probabilities $\Pr(t(k)=i)=\frac{1}{M_{i}A_{k}}$ , then ${\mathbb{E}}\left[\nabla F(\mathbf{x}_{t(k)})\right]=\frac{1}{A_{k}}\sum_{i=0}^{k-1}\frac{G_{i+1}^{2}}{M_{i}}$. ∎
@@ Appendix C Missing Proofs in Section 6
Lemma 9.
Under Assumptions 1, 3, 4 and 5, for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, if $\gamma_{k}\coloneqq\frac{3\sqrt{2}L}{M_{k}-L}$, $a\coloneqq\frac{\sqrt{2}}{8L}$, and $M_{k}>(1+3\sqrt{2})L$, we have:
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{48M_{k}}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]-\frac{3\sqrt{2}}{2M_{k}}\Delta_{k}+\frac{27\sqrt{2}L\sigma^{2}}{4M_{k}^{2}}.$
@@
We repeat the proof of Lemma 5 but now require that $\frac{L^{2}}{\gamma_{k}}\bigl{(}\frac{8}{(M_{k}-L)^{2}}+\frac{1}{M_{k}^{2}+L^{2}}\bigr{)}\leq\frac{\gamma_{k}}{2}$. To satisfy this inequality, it suffices to choose $\gamma_{k}=\frac{3\sqrt{2}L}{M_{k}-L}$, which requires that $M_{k}>(1+3\sqrt{2})L$. Now we have $a=\frac{\sqrt{2}}{8L}$ and:
$\Phi_{k+1}\leq\Phi_{k}-\frac{M_{k}-L}{24(M_{k}^{2}+L^{2})}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]-\frac{3\sqrt{2}L}{2(M_{k}-L)}\Delta_{k}+\frac{27\sqrt{2}L\sigma^{2}}{4(M_{k}-L)^{2}}.$
Let $M_{k}=\tau_{k}L$ for some $\tau_{k}>(1+3\sqrt{2})$. Note that, for $\tau_{k}>(1+3\sqrt{2})$, we have $\frac{\tau_{k}-1}{\tau_{k}^{2}+1}\geq\frac{1}{2\tau_{k}}$, $\frac{1}{2(\tau_{k}-1)}\geq\frac{1}{2\tau_{k}}$ and $\frac{1}{(\tau_{k}-1)^{2}}\leq\frac{2}{\tau_{k}^{2}}$. Therefore, for $M_{k}>(1+3\sqrt{2})L$:
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{48M_{k}}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]-\frac{3\sqrt{2}}{2M_{k}}\Delta_{k}+\frac{27\sqrt{2}L\sigma^{2}}{4M_{k}^{2}}.\qed$
Appendix D Missing Proofs in Section 7
First, we notice that, under Assumptions 1 and 3, lemma 2 still holds in the inexact case. Now we give an analogous result to Lemma 4 in the inexact case:
Lemma 10.
Under Assumption 3, for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, we have:
$(M_{k}^{2}+L^{2})R_{k}\geq\frac{1}{4}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]-\Delta_{k}-{\mathbb{E}}\left[\left\lVert\nabla\Omega(\mathbf{x}_{k+1})\right\rVert^{2}\right].$
Proof.
Indeed,
@@ Appendix D Missing Proofs in Section 7
$=\left\lVert M_{k}(\mathbf{x}_{k}-\mathbf{x}_{k+1})+\nabla\Omega(\mathbf{x}_{k+1})+(\nabla f(\mathbf{x}_{k})-\mathbf{m}_{k})+(\nabla f(\mathbf{x}_{k+1})-\nabla f(\mathbf{x}_{k}))\right\rVert^{2}$
$\leq 4M_{k}^{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+4\left\lVert\nabla\Omega(\mathbf{x}_{k+1})\right\rVert^{2}+4\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}+4\left\lVert\nabla f(\mathbf{x}_{k+1})-\nabla f(\mathbf{x}_{k})\right\rVert^{2}$
$\leq 4(M_{k}^{2}+L^{2})\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+4\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}+4\left\lVert\nabla\Omega(\mathbf{x}_{k+1})\right\rVert^{2}.$
Taking expectations and rearranging, we obtain the claim. ∎
Notice that Lemma 1 still holds. We continue to use the same Lyapunov function $\Phi_{k}\coloneqq F_{k}+a\Delta_{k}$.
Lemma 11.
Under Assumptions 1, 3 and 4, for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, if $\gamma_{k}\coloneqq\sqrt{\frac{456}{17}}\frac{L}{M_{k}-L}$, $a\coloneqq\sqrt{\frac{19}{408}}\frac{1}{L}$, that in each iteration condition (10) holds. Then, for $M_{k}>7L$, we have:
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{68M_{k}}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{12L\sigma^{2}}{M_{k}^{2}}+\frac{2S_{k}}{17M_{k}}.$ (11)
@@
Denote $G_{k+1}^{2}\coloneqq{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]$. Plugging Equation 10 into Lemma 10, we obtain
$(M_{k}^{2}+L^{2})R_{k}\geq\frac{1}{4}G_{k+1}^{2}-\Delta_{k}-{\mathbb{E}}\left[\left\lVert\nabla\Omega(\mathbf{x}_{k+1})\right\rVert^{2}\right]\geq\frac{1}{4}G_{k+1}^{2}-\Delta_{k}-\frac{M_{k}^{2}}{16}R_{k}-S_{k}.$